### v1

@@ 1 Introduction
Power transmission corridor inspection is crucial for maintaining the electric power transmission safety and stability [13, 10, 18]. Current inspection methods based on object detection have achieved great progress in locating hazard targets around transmission corridors [26, 59]. However, they typically predict rough bounding boxes, hindering subsequent tasks (morphology analysis, ranging) that require accurate hazard shape. Therefore, power transmission corridor hazard segmentation (PTCHS) is designed to further separate transmission equipment and surrounding hazards from complex background, providing a more practical solution for the following risk assessment and prevention strategy development. Related studies have achieved promising results [10, 60]. Nevertheless, the lack of precise segmentation limits the practical effectiveness of PTCHS.
@@
Recently, for general image segmentation, the Segment Anything Model (SAM) [20] has emerged as a foundational model. Benefiting from large-scale pretraining, SAM demonstrates impressive transferability and adaptation across numerous tasks in diverse scenarios, including matting [23, 52], medical image segmentation [9, 47], and hierarchical text segmentation [54]. Motivated by the excellent properties of SAM, we leverage it to advance the PTCHS task. However, there are two challenges. 1) Data scarcity. There are few PTCHS dataset available, lacking adequate hazard categories and high-quality mask annotations in complex transmission corridor scenarios. 2) Segmentation quality. Even after dedicated fine-tuning on PTCHS data, SAM struggles to deal with target objects with fine structure in complex background. For instance, as shown in the second and fourth columns of Fig. 1, the segmentation on tower crane from SAM is severely interfered by surrounding buildings or equipment.
To solve the above challenges, we start by constructing the first large-scale benchmark for PTCHS, ELE-40K, containing 44,094 pixel-level annotated image-mask pairs derived from real-world transmission corridors. ELE-40K covers transmission equipment, construction vehicles, and environmental hazards, such as wildfire, smoke, etc. To obtain the mask annotations, we adopt a semi-automatic and iterative strategy [20]. We first fine-tune SAM with some manual labeled data, and then annotate the remaining images incorporating the iteratively retrained SAM with manual rectification. Moreover, we propose ELE-SAM for advancing SAM in PTCHS task. Technically, we introduce an Context-Aware Prompt Adapter to improve the prompt tokens. Besides, to better deal with the hazard objects with fine structure during mask decoding, we design a High-Fidelity Mask Decoder. In the mask decoder, multi-sourced features are fused, iteratively refined, and then scaled to a higher resolution of $512\times 512$. Experimental results demonstrate the effectiveness of our method on the hard samples which require high-quality segmentation capability.
@@
The paper is organized as follows: a brief review of related works is presented in Sec. 2. The proposed method and dataset are illustrated in Sec. 3. Extensive experiments are reported in Sec. 4. The paper finally concludes in Sec. 5.
@@
NAL-RGB [39] 3,568 Binary Mask No 1 -
PLDU [62] 573 Binary Mask Yes 1 -
PLDM [62] 287 Binary Mask Yes 1 -
Vepl [5] 3,724 Binary Mask Yes 2 1
@@
SR-RGB [51] 2,000 Class Label Yes 1 -
TTPLA [1] 1,100 Binary Mask Yes 4 -
@@ 2.1 Power Transmission Corridor Inspection
Recently, the application of computer vision technology in power transmission corridor inspection arouses increasing attention [44, 33]. Object detection methods are primarily adopted. These methods are categorized into two-stage and single-stage algorithms. Two-stage methods extract candidate regions before detection, offering higher accuracy [36, 46], while single-stage methods perform end-to-end detection using convolutional neural networks, prioritizing speed [59, 45]. However, these object detection methods solely predict bounding boxes that include background objects, hindering precise contour delineation of hazardous objects. This limitation impedes morphological analysis and reduces the accuracy of tasks like distance measurement. In contrast, power transmission corridor hazard segmentation (PTCHS) emerges as an optimal approach [40, 35, 19]. Wei et al. [49] reduce the cost of manual labeling and improve the performance of power line segmentation based on the Swin-Unet framework with improved linear embedding and efficient sample synthesis techniques. Similarly, Hu et al. [19] introduce a gated axial attention mechanism and a local normalisation module for axial channels based on UNet++ [58], which effectively extend the receptive field and capture the relative position deviation. Abdelfattahet al. [2] propose a novel GAN [17] framework, which leverages adversarial training, a Hough transform loss function, and a semantic decoder to achieve excellent performance in segmenting power lines. However, these methods are limited to specific transmission equipment or hazards and cannot address dynamic risks in real-world transmission corridors. Thus, a generalized segmentation model with robust adaptability is urgently required to address this challenge.
@@ 2.2 Segment Anything Model
The Segment Anything Model (SAM) [20] advances the segmentation field through large-scale pretraining and enabling interactive visual prompts. The extraordinary performance and favorable generalization across various real-world scenarios render SAM as one of the vision foundation models. However, SAM’s zero-shot performance in a few specialized fields exhibits a notable decrease when encountering unseen features [48, 29]. To promote the adaptability, researchers have improved the architecture of SAM to better deal with specialized challenges, such as camouflaged object detect [7] and medical image analysis  [8, 30, 14]. For instance, Chen et al. [7] enhance SAM’s prompt adaptation capability by introducing a response filter and semantic matcher, thereby improving mask quality in camouflaged object detection scenarios. Cheng et al. [8] employ hierarchical decoding with mask-guided self-attention to enhance SAM’s adaptation for medical image segmentation in the context of limited data. In addition, HQ-SAM [21] enhances SAM’s segmentation precision by integrating a learnable High-Quality Output Token within the mask decoder, leveraging effective fusion of features from ViT layers. PA-SAM [50] refines SAM’s segmentation capabilities through a prompt adapter, which optimizes feature extraction and decoding for improved flexibility and performance. Hi-SAM [54], specifically tailored for text stroke segmentation, employs a hierarchical mask decoder to fine-tune SAM, enabling unified segmentation and layout analysis, thereby excelling in text-centric applications. In contrast, to overcome the limitations of SAM in the PTCHS task, we introduce ELE-SAM, a tailored framework for PTCHS. ELE-SAM integrates a Context-Aware Prompt Adapter, a High-Fidelity Mask Decoder, and leverages the ELE-40K dataset to address critical challenges, including fine-structure hazard detection, complex background segmentation, and data scarcity.
Figure 2: The overall architecture of ELE-SAM. ELE-SAM employs an encoder-decoder paradigm as SAM, incorporating two novel modules, the context-aware prompt adapter (CAPA) and the high-fidelity mask decoder (HFMD). CAPA generates more discriminative prompt tokens to better distinguish the target object. To segment the target object in high-quality, HFMD further produces mask features in a higher resolution of $512\times 512$, overcoming the perception loss on object details with the original mask feature resolution.
@@ 2.3 Benchmark Datasets
Publicly available datasets for PTCHS are scarce, limiting the development and evaluation of advanced detection models. General datasets like COCO [42] and ImageNet [11] lack power-specific annotations, focusing on generic objects rather than the structurally complex hazards and transmission equipment. While some small-scale datasets capture power transmission scenes via drone or high-resolution imagery, their limited scope and variety hinder model generalization across diverse settings [5, 28, 6, 1, 32]. For instance, as presented in Tab. 1, although the TTPLA [1] introduces mask annotations, it primarily focuses on a limited range of transmission equipment, such as utility poles and lines, while critical components like transmission towers remain sparsely annotated. Furthermore, the dataset lacks attention to potential hazards in the surrounding environment that could pose threats to transmission equipment. These limitations collectively render existing datasets insufficient for providing comprehensive training in the context of power transmission corridor inspection. To this end, we introduce the ELE-40K dataset, designed to include diverse equipment and hazard scenarios, providing a robust benchmark to improve model accuracy and reliability in real-world PTCHS.
@@ 3.1 Overview of ELE-SAM
As depicted in Fig. 2, ELE-SAM consists of four major components: 1) a frozen image encoder from SAM [20], 2) a frozen SAM’s prompt encoder for encoding the initial prompt tokens, 3) a plug-and-play Context-Aware Prompt Adapter (CAPA) for mining more discriminative prompt tokens, and 4) a customized High-Fidelity Mask Decoder (HFMD) for producing and refining mask features in bi-resolution.
Concretely, given the input image $\bm{I}$, the image encoder generates image embedding $\bm{I}_{emb}$. Following [21], we also extract and fuse the early layer and final layer features from image encoder, resulting in the fused features $\bm{I}_{fusion}$. Note that the image encoder is the same as that in SAM, without any learnable component inserted. Meanwhile, visual prompts like boxes are embedded by the prompt encoder and combined with output token, forming the initial prompt tokens $\bm{P}$.
In HFMD, the two-way Transformer decoder layers firstly interact with CAPA to generates enhanced prompt tokens, including quantity-augmented sparse prompt tokens. Specifically, with the image $\bm{I}$, image embedding $\bm{I}_{emb}$, and prompt tokens $\bm{P}$, after the two-way decoder layers that interact with CAPA, promoted image embedding $\bm{I}_{emb}^{\prime}$ and enhanced prompt tokens $\bm{P}^{\prime}$ are obtained. Then, $\bm{I}_{emb}^{\prime}$ is upsampled to $256\times 256$ in resolution, forming the mask features $\bm{F}$. After the final token-to-image attention, the output token $\bm{T}$ is separated from $\bm{P}^{\prime}$ for subsequent mask prediction. Finally, in our newly introduced modules for producing and refining mask features in bi-resolution, given fused features $\bm{I}_{fusion}$, mask features $\bm{F}$, and the output token $\bm{T}$, mask logits with $256\times 256$ and $512\times 512$ resolution can be achieved respectively.
In the following subsections, we delve into the technical details of CAPA and HFMD.
@@ 3.2 Context-Aware Prompt Adapter
Given the initial prompt tokens $\bm{P}$, $\bm{P}$ could be coarse and inadequate to determine some uncertain regions in high-quality segmentation. Inspired by PA-SAM [50], we incorporate its Prompt Adapter (PA) to generate more distinctive prompt tokens. Differently, we additionally design a Local Context Enhancer (LCE) and insert it before PA to enable pre-interaction between prompts, thereby enhancing the adaptive prompt generation. The structure of LCE is illustrated in Fig. 3.
@@
The prompt adapter $PA$ consists of two key components: 1) Adaptive Detail Enhancement. This scheme explores detail information from the image and its Canny gradient by Dense Prompt Compensation and Sparse Prompt Optimization. 2) Hard Point Mining. This operation samples more positive and negative points, then embeds and concatenates them to the input prompts $\hat{\bm{P}}$. In the prompt adapter, coarse, refined, and uncertain masks are produced and used to calculate the loss $\mathcal{L}_{\text{PA}}$. We follow the same implementation as in PA-SAM [50], where more details are expanded.
@@ 3.3 High-Fidelity Mask Decoder
Figure 5: As illustrated in the figure, the data distribution of our ELE-40K dataset is clearly illustrated. Encompassing various transmission equipment and surrounding hazards under conditions such as clear skies, overcast, fog, rain, and darkness. This highlights the comprehensiveness of the ELE-40K dataset.
@@
Figure 7: From the first to the third row, we present the ground truth, the suboptimal method (HQ-SAM [21]), and our ELE-SAM, respectively. Specifically, (a) illustrates the visualization results on ELE-40K dataset, while (b) shows the results on HQSeg44K dataset.
@@ 3.4 Loss Function
Since ELE-SAM predicts bi-resolution results, we apply supervision for both of them to ensure coarse-to-fine mask evolution. In particular, given low-resolution prediction $\bm{M}_{\text{lr}}$ and high-resolution prediction $\bm{M}_{\text{hr}}$, we calculate the Dice loss [31] for each resolution level with ground-truth $\bm{M}_{\text{GT}}$. The total loss $\mathcal{L}$ equals to the Dice loss from two resolution levels plus the loss $\mathcal{L}_{\text{PA}}$ from prompt adapter, which is formulated as:
@@ 3.5 ELE-40K Benchmark
Dataset Sources. ELE-40K dataset is derived from two primary sources: 1) Self-collected dataset: We collected and annotated a total of 22,878 image-mask pairs captured by unmanned aerial vehicles and cameras along transmission corridors. 2) Publicly available datasets: We extracted and re-annotated data from existing publicly available datasets  [1, 12, 15, 27, 38], covering instances such as power lines, insulators, poles and engineering vehicles. Combining these two sources, we obtained a total of 31,205 annotated transmission equipment and 12,889 hazards image-mask pairs.
Dataset Construction and Statistic. The data from public image repositories in ELE-40K are real-world transmission corridor images, capturing diverse operational and environmental conditions. To achieve high-quality annotations, inspired by SAM [20], we employ a semi-automated annotation process. The process includes: 1) Manual initial annotating. Domain experts annotate images at the pixel level, identifying transmission equipment and potential hazards. In the initial stage, we annotated 1,000 image-mask pairs. 2) We train ELE-SAM with pixel-level annotation from step 1. The derived model is then applied to predict initial masks for a new set of 2,000 images, accelerating the annotation process. 3) Experts manually refine the false positive and false negative regions in the preliminary masks generated in step 2 to ensure annotation accuracy. 4) We combine the data in step 1 and refined data for retraining ELE-SAM to improve accuracy in subsequent iterations. The loop of SAM-assisted labeling, manual refinement, and retraining is repeated until all data is fully annotated. Finally, the dataset contains key transmission equipment such as transmission towers, poles, and support structures. Frequent hazards encountered during maintenance activities are also included, such as engineering vehicles, bulldozers, cranes, and other heavy machinery. To further enhance applicability, high-impact environmental hazards, such as wildfire and smoke, are also included. ELE-40K features different imaging condition, including froggy, rainy, and darkness. Some examples are shown in Fig. 5. The detailed instance distribution is illustrated in Fig. 6.

### v2

@@ 1 Introduction
Power transmission corridor inspection is crucial for maintaining the electric power transmission safety and stability [9802679, 9699431, 8972612]. Current inspection methods based on object detection have achieved great progress in locating hazard targets around transmission corridors [LI2024123983, 9530535]. However, they typically predict rough bounding boxes, hindering subsequent tasks (morphology analysis, ranging) that require accurate hazard shape. Therefore, power transmission corridor hazard segmentation (PTCHS) aims to further separate transmission equipment and surrounding hazards from complex background, providing a more practical solution for the following risk assessment and prevention strategy development. While PTCHS does not introduce a fundamentally new segmentation paradigm, it addresses the critical real-world need for hazard screening in power transmission corridors. Related studies have achieved promising results [9699431, Zhou2022DualViewSA]. Nevertheless, the lack of precise segmentation limits the practical effectiveness of PTCHS.
@@
Recently, for general image segmentation, the Segment Anything Model (SAM) [kirillov2023segment] has emerged as a foundational model. Benefiting from large-scale pretraining, SAM demonstrates impressive transferability and adaptation across numerous tasks in diverse scenarios, including matting [li2024matting], medical image segmentation [cheng2023sam], and hierarchical text segmentation [hisam]. Motivated by the excellent properties of SAM, we leverage it to advance the PTCHS task. However, there are two challenges. 1) Data scarcity. There are few PTCHS dataset available, lacking adequate hazard categories and high-quality mask annotations in complex transmission corridor scenarios. 2) Segmentation quality. Even after dedicated fine-tuning on PTCHS data, SAM struggles to deal with target objects with fine structure in complex background. For instance, as shown in the second and fourth columns of Fig. 1, the segmentation on tower crane from SAM is severely interfered by surrounding buildings or equipment.
To solve the above challenges, we start by constructing the first large-scale benchmark for PTCHS, ELE-40K, containing 44,094 pixel-level annotated image-mask pairs derived from real-world transmission corridors. ELE-40K covers transmission equipment, construction vehicles, and environmental hazards, such as wildfire, smoke, etc. To obtain the mask annotations, we adopt a semi-automatic and iterative strategy (Kirillov et al., 2023). We first fine-tune SAM with some manual labeled data, and then annotate the remaining images incorporating the iteratively retrained SAM with manual rectification. For PTCHS on ELE-40K, we propose a baseline model named ELE-SAM. Technically, we introduce an Context-Aware Prompt Adapter to improve the prompt tokens. Besides, to better deal with the hazard objects with fine structure during mask decoding, we design a High-Fidelity Mask Decoder. In the mask decoder, multi-sourced features are fused, iteratively refined, and then scaled to a higher resolution of $512\times 512$. Experimental results demonstrate the effectiveness of our method on the hard samples which require high-quality segmentation capability.
@@
The paper is organized as follows: a brief review of related works is presented in Sec. 2. The proposed method and dataset are illustrated in Sec. 3. Extensive experiments are reported in Sec. 4. Some limitations are discussed in Sec. 5. The paper finally concludes in Sec. 6.
@@
NAL-RGB [inbook] 3,568 Binary Mask No 1 -
PLDU [rs11111342] 573 Binary Mask Yes 1 -
PLDM [rs11111342] 287 Binary Mask Yes 1 -
Vepl [data8080128] 3,724 Binary Mask Yes 2 1
@@
SR-RGB [Yetgin2017PowerlineID] 2,000 Class Label Yes 1 -
TTPLA [abdelfattah2020ttplaaerialimagedatasetdetection] 1,100 Binary Mask Yes 4 -
@@ 2.1 Power Transmission Corridor Inspection
Recently, the application of computer vision technology in power transmission corridor inspection arouses increasing attention [10445528, Nair2024]. Object detection methods are primarily adopted. These methods are categorized into two-stage and single-stage algorithms. Two-stage methods extract candidate regions before detection, offering higher accuracy [9301267, https://doi.org/10.1049/gtd2.12688], while single-stage methods perform end-to-end detection using convolutional neural networks, prioritizing speed [9530535, Tang2024]. However, these object detection methods solely predict bounding boxes that include background objects, hindering precise contour delineation of hazardous objects. This limitation impedes morphological analysis and reduces the accuracy of tasks like distance measurement. In contrast, power transmission corridor hazard segmentation (PTCHS) emerges as an optimal approach [SHEN2023103263, DBLP:journals/corr/RonnebergerFB15, rs16234585]. Wei et al[Wei2022] reduce the cost of manual labeling and improve the performance of power line segmentation based on the Swin-Unet framework with improved linear embedding and efficient sample synthesis techniques. Hu et al[rs16234585] introduce a gated axial attention mechanism and a local normalization module for axial channels. Abdelfattahet al[10301645] propose a novel framework, which leverages adversarial training, a Hough transform loss function, and a semantic decoder to achieve excellent performance in segmenting power lines. However, these methods are limited to specific transmission equipment or hazards and cannot address dynamic risks in real-world transmission corridors. Thus, a generalized segmentation model with robust adaptability is urgently required to address this challenge.
@@ 2.2 Segment Anything Model
The Segment Anything Model (SAM) [kirillov2023segment] advances the segmentation field through large-scale pretraining and enabling interactive visual prompts. The extraordinary performance and favorable generalization across various real-world scenarios render SAM as one of the vision foundation models. However, SAM’s zero-shot performance in a few specialized fields exhibits a notable decrease when encountering unseen features [wang2023sammeetsroboticsurgery, mohapatra2023samvsbetcomparative]. To promote the adaptability, researchers have improved the architecture of SAM to better deal with specialized challenges, such as camouflaged object detect [chen2024samcodsamguidedunifiedframework] and medical image analysis  [Ma2024, 10656025]. For instance, Chen et al[chen2024samcodsamguidedunifiedframework] enhance SAM’s prompt adaptation capability by introducing a response filter and semantic matcher, thereby improving mask quality in camouflaged object detection scenarios. HQ-SAM [sam_hq] enhances SAM’s segmentation precision by integrating a learnable High-Quality Output Token within the mask decoder, leveraging effective fusion of features from ViT layers. PA-SAM [xie2024pasampromptadaptersam] refines SAM’s segmentation capabilities through a prompt adapter, which optimizes feature extraction and decoding for improved flexibility and performance. Hi-SAM [hisam] realizes hierarchical text segmentation in a unified framework. In contrast, to overcome the limitations of SAM in the PTCHS task, we introduce ELE-SAM, a tailored framework for PTCHS. ELE-SAM integrates a Context-Aware Prompt Adapter, a High-Fidelity Mask Decoder, and leverages the ELE-40K dataset to address critical challenges, including fine-structure hazard detection, complex background segmentation, and data scarcity.
Figure 2: The overall architecture of ELE-SAM. ELE-SAM employs an encoder-decoder paradigm as SAM, incorporating two novel modules, the context-aware prompt adapter (CAPA) and the high-fidelity mask decoder (HFMD). CAPA generates more discriminative prompt tokens to better distinguish the target object. For high-quality segmentation, HFMD further produces mask features in a higher resolution of $512\times 512$, overcoming the perception loss on object details with the original mask feature resolution.
@@ 2.3 Benchmark Datasets
Publicly available datasets for PTCHS are scarce, limiting the development and evaluation of advanced detection models. General datasets like COCO [singh2024benchmarkingobjectdetectorscoco] and ImageNet [5206848] lack power-specific annotations, focusing on generic objects rather than the structurally complex hazards and transmission equipment. While some small-scale datasets capture power transmission scenes via drone or high-resolution imagery, their limited scope and variety hinder model generalization across diverse settings [data8080128, abdelfattah2020ttplaaerialimagedatasetdetection, 8610301]. For instance, as presented in Tab. 1, although the TTPLA [abdelfattah2020ttplaaerialimagedatasetdetection] introduces mask annotations, it primarily focuses on a limited range of transmission equipment, such as utility poles and lines, while critical components like transmission towers remain sparsely annotated. Furthermore, the dataset lacks attention to potential hazards in the surrounding environment that could pose threats to transmission equipment. These limitations collectively render existing datasets insufficient for providing comprehensive training in the context of power transmission corridor inspection. To this end, we introduce the ELE-40K dataset, designed to include diverse equipment and hazard scenarios, providing a robust benchmark to improve model accuracy and reliability in real-world PTCHS.
@@ 3.1 Overview of ELE-SAM
As depicted in Fig. 2, ELE-SAM consists of four major components: 1) a frozen image encoder from SAM [kirillov2023segment], 2) a frozen SAM’s prompt encoder for encoding the initial prompt tokens, 3) a plug-and-play Context-Aware Prompt Adapter (CAPA) for mining more discriminative prompt tokens, and 4) a customized High-Fidelity Mask Decoder (HFMD) for producing and refining mask features in bi-resolution.
Concretely, given the input image $\bm{I}$, the image encoder generates image embedding $\bm{I}_{emb}$. Following [sam_hq], we also extract and fuse the early layer and final layer features from image encoder, resulting in the fused features $\bm{I}_{fusion}$. Note that the image encoder is the same as that in SAM, without any learnable component inserted. Meanwhile, visual prompts like boxes are embedded by the prompt encoder and combined with output token, forming the initial prompt tokens $\bm{P}$.
In HFMD, the two-way Transformer decoder layers firstly interact with CAPA to generates enhanced prompt tokens, including quantity-augmented sparse prompt tokens. Specifically, with the image $\bm{I}$, image embedding $\bm{I}_{emb}$, and prompt tokens $\bm{P}$, after the two-way decoder layers that interact with CAPA, promoted image embedding $\bm{I}_{emb}^{\prime}$ and enhanced prompt tokens $\bm{P}^{\prime}$ are obtained. Then, $\bm{I}_{emb}^{\prime}$ is upsampled to $256\times 256$ in resolution, forming the mask features $\bm{F}$. After the final token-to-image attention, the output token $\bm{T}$ is separated from $\bm{P}^{\prime}$ for subsequent mask prediction. Finally, in our newly introduced modules for producing and refining mask features in bi-resolution, given fused features $\bm{I}_{fusion}$, mask features $\bm{F}$, and the output token $\bm{T}$, mask logits with $256\times 256$ and $512\times 512$ resolution can be achieved respectively. In the following subsections, we delve into the technical details of CAPA and HFMD.
@@ 3.2 Context-Aware Prompt Adapter
Given the initial prompt tokens $\bm{P}$, $\bm{P}$ could be coarse and inadequate to determine some uncertain regions in high-quality segmentation. Inspired by PA-SAM [xie2024pasampromptadaptersam], we incorporate its Prompt Adapter (PA) to generate more distinctive prompt tokens. Differently, we additionally design a Local Context Enhancer (LCE) and insert it before PA to enable pre-interaction between prompts, thereby enhancing the adaptive prompt generation. The structure of LCE is illustrated in Fig. 3.
@@
The prompt adapter $PA$ consists of two key components: 1) Adaptive Detail Enhancement. This scheme explores detail information from the image and its Canny gradient by Dense Prompt Compensation and Sparse Prompt Optimization. 2) Hard Point Mining. This operation samples more positive and negative points, then embeds and concatenates them to the input prompts $\hat{\bm{P}}$. In the prompt adapter, coarse, refined, and uncertain masks are produced and used to calculate the loss $\mathcal{L}_{\text{PA}}$. We follow the same implementation as in PA-SAM [xie2024pasampromptadaptersam], where more details are expanded.
@@ 3.3 High-Fidelity Mask Decoder
Figure 5: Samples in ELE-40K under different imaging conditions. It shows transmission equipment and surrounding hazards under various conditions such as sunny, overcast, fog, rain, and darkness. This highlights the comprehensiveness of the dataset.
@@
((a))
((b))
Figure 7: From the first to the third row, we present the ground truth, the suboptimal method (HQ-SAM [sam_hq]), and our ELE-SAM, respectively. Specifically, (a) illustrates the visualization results on ELE-40K dataset, while (b) shows the results on HQSeg44K dataset.
@@ 3.4 Loss Function
Since ELE-SAM predicts bi-resolution results, we apply supervision for both of them to ensure coarse-to-fine mask evolution. In particular, given low-resolution prediction $\bm{M}_{\text{lr}}$ and high-resolution prediction $\bm{M}_{\text{hr}}$, we calculate the Dice loss [dice] for each resolution level with ground-truth $\bm{M}_{\text{GT}}$. The total loss $\mathcal{L}$ equals to the Dice loss from two resolution levels plus the loss $\mathcal{L}_{\text{PA}}$ from prompt adapter, which is formulated as:
@@ 3.5 ELE-40K Benchmark
Dataset Sources. ELE-40K dataset is derived from two primary sources: 1) Self-collected dataset: We collected and annotated a total of 22,878 image-mask pairs captured by unmanned aerial vehicles and cameras along transmission corridors. 2) Publicly available datasets: We extracted and re-annotated data from existing publicly available datasets  [abdelfattah2020ttplaaerialimagedatasetdetection, towerdetection_dataset, electricpolefull-yxgbo_dataset, cranes-xxz5d_dataset, excavators-cwlh0_dataset], covering instances such as power lines, insulators, poles and engineering vehicles. Combining these two sources, we obtained a total of 31,205 annotated transmission equipment and 12,889 hazards image-mask pairs.
Dataset Construction and Statistic. The images in ELE-40K are from real-world transmission corridor scenario, capturing diverse operational and environmental conditions. To achieve segmentation annotations, inspired by SAM (Kirillov et al., 2023), we employ a semi-automated annotation process. The process includes: 1) Manual initial annotating. Domain experts annotate images at the pixel level, identifying transmission equipment and potential hazards. In the initial stage, we annotated 1,000 image-mask pairs. 2) We train ELE-SAM with pixel-level annotation from step 1. The derived model is then applied to predict initial masks for a new set of 2,000 images, accelerating the annotation process. 3) Experts manually refine the false positive and false negative regions in the preliminary masks generated in step 2 to ensure annotation accuracy. 4) We combine the data in step 1 and refined data for retraining ELE-SAM to improve accuracy in subsequent iterations. The loop of SAM-assisted labeling, manual refinement, and retraining is repeated until all data is fully annotated.
Due to annotation cost and the practical requirement that subsequent tasks like distance measurement and risk assessment rely on overall object contours, the hollows between steel wires inside the objects have not been accurately annotated. During evaluation, we follow the promptable segmentation paradigm as in HQSeg-44K, where each object is individually segmented based on the given bounding box prompt. As modern detectors can provide high-accuracy boxes in real-time, we focus on category-agnostic promptable segmentation. Finally, the dataset contains key transmission equipment such as transmission towers, poles, and support structures. Frequent hazards encountered during maintenance activities are also included, such as engineering vehicles, bulldozers, cranes, and other heavy machinery. To further enhance applicability, high-impact environmental hazards, such as wildfire and smoke, are also included. ELE-40K features different imaging condition, including froggy, rainy, and darkness. Some examples are shown in Fig. 5. The detailed instance distribution is illustrated in Fig. 6.